OpenStudy (anonymous):

The airliner concorde flies 2000km at a speed of 1600km/hr and then returns due to bad weather at a speed of 1000km/h find the average speed for the whole trip , I solved it bt not sure of my answer help pls?

OpenStudy (anonymous):

That would be 1300km/h, simple if the plane goes any distance; and the speed was 1600 and on the return trip of the same distance; you need to find the average speed b/w them

OpenStudy (anonymous):

Can u solve it if you dont mind?

OpenStudy (mathstudent55):

1300km/h is not correct. The average speed is not the average of the averages. An average speed is calculated by dividing the total distance traveled by the total time it took to travel. What is the total distance traveled? What is the total time of travel?

OpenStudy (anonymous):

Total distance travelled is 4000km and total time is totaldistance/totalspeed so 4000/2600?

OpenStudy (mathstudent55):

The total distance is correct, 4000 km. The total time of travel is made up of two parts: 2000 km at 1600 km/h and 2000 km at 1000 km/h. Find the time of each part and add the times together.

OpenStudy (mathstudent55):

Averages can't be added. There is no total average speed of 2600 km/h. Each part of the trip has to be calculated on its own.

OpenStudy (anonymous):

I got the answer , thanks alot :D

OpenStudy (mathstudent55):

What did you get?

OpenStudy (anonymous):

3.25km/h

OpenStudy (mathstudent55):

You're on the way to the answer. Remember, any problem you solve needs to get an answer that makes sense. The speeds were 1600 km/h and 1000 km/h. An overall average speed of 3.25 km/h makes no sense. 3.25 h is the total travel time. 2000 km / (1600 km/h) = 1.25 h (amount of time going) 2000 km / (1000 km/h) = 2 h (amount of time returning) 1.25 h + 2 h = 3.25 h (total travel time) Now divide 4000 km by 3.25 h to find the overall average speed.

OpenStudy (anonymous):

Yeach I actually realized it made no sense 4000÷3.25=1230.8 (3 s.f)

OpenStudy (mathstudent55):

Now you got it. Great job!

OpenStudy (anonymous):

Thanks again ^.^

OpenStudy (mathstudent55):

You mean 3 significant figures? If so, then the answer is 1230 km/h